# MagicLens[Uncaptioned image]: Self-Supervised Image Retrieval with Open-Ended Instructions

### v1

@@ Abstract
Image retrieval, i.e., finding desired images given a reference image, inherently encompasses rich, multi-faceted search intents that are difficult to capture solely using image-based measures. Recent work leverages text instructions to allow users to more freely express their search intents. However, existing work primarily focuses on image pairs that are visually similar and/or can be characterized by a small set of pre-defined relations. The core thesis of this paper is that text instructions can enable retrieving images with richer relations beyond visual similarity. To show this, we introduce MagicLens, a series of self-supervised image retrieval models that support open-ended instructions. MagicLens is built on a key novel insight: image pairs that naturally occur on the same web pages contain a wide range of implicit relations (e.g., inside view of), and we can bring those implicit relations explicit by synthesizing instructions via large multimodal models (LMMs) and large language models (LLMs). Trained on 36.7M (query image, instruction, target image) triplets with rich semantic relations mined from the web, MagicLens achieves comparable or better results on eight benchmarks of various image retrieval tasks than prior state-of-the-art (SOTA) methods. Remarkably, it outperforms previous SOTA but with a 50$\times$ smaller model size on multiple benchmarks. Additional human analyses on a 1.4M-image unseen corpus further demonstrate the diversity of search intents supported by MagicLens.
@@
$1$Google DeepMind    $2$The Ohio State University
https://open-vision-language.github.io/MagicLens/
[…]
@@ 1 Introduction
Figure 1: Top-1 retrieved images using MagicLens and the prior SOTA (Gu et al., 2024) from a retrieval pool containing 1.4M images. The prior SOTA method, while accepting text instructions, primarily retrieves images based on visual similarity to the query image, ignoring the nuances of the text instructions. In contrast, MagicLens excels at retrieving both visually similar images and those that align with the deeper meaning and context of the text instructions — even when the images do not resemble the query. For example, if given a query image of the Burj Al Arab and the instruction “find other attractions in this country”, it can successfully locate images of the Palm Islands in Dubai.
@@
In this paper, we present MagicLens, a series of self-supervised image retrieval models trained on a wide range of (query image, instruction, target image) triplets that reflect naturally occurring semantic relations, mined from web pages and curated with state-of-the-art (SOTA) foundation models. Specifically, we extract image pairs that naturally occur on the same web page to form positive pairs that carry abundant but natural semantic relations. We then apply both large multimodal models (Chen et al., 2023b, a) and large language models (Anil et al., 2023) to refine the description of such open-ended semantic relation, into open-ended instruction. Figure 2 presents an overview of the data construction pipeline. As a concrete example, a camera review website [^1] https://amateurphotographer.com/review/nikon-z5-review/ that presents the image of a Nikon Camera and the image of a Nikon Charger would offer an interesting and non-trivial relation “charger of a product”, which would then be curated by the LMM$+$LLM pipeline, and produce a final instruction Find a charger for it. This process produces open-ended instructions that depict diverse semantic relations beyond mere visual similarity, resulting in a large-scale training dataset with 36.7M high-quality triplets over a wide distribution.
With the constructed dataset, we train dual-encoder models called MagicLens, which retrieve images given a query consisting of an image with an instruction. Our models achieve comparable or better results than prior SOTA methods on eight benchmarks, ranging from various multimodality-to-image and image-to-image retrieval tasks. In addition, MagicLens can retain or even significantly improve the text-to-image retrieval performance of the underlying single-modality encoders. With a 50 times smaller model size than prior SOTA methods, MagicLens outperforms them on multiple benchmarks: CIRCO (Baldrati et al., 2023), Domain Transfer ImageNet (Saito et al., 2023), and GeneCIS (Vaze et al., 2023). To further examine our models’ capabilities in a more realistic scenario, we construct the largest retrieval pool to date with 1.4 million unseen images and perform retrieval given human-written search queries with diverse instructions. The human evaluation finds that MagicLens can successfully satisfy complex and beyond visual search intents, whereas the prior SOTA fails to do so.
Figure 3: Data construction pipeline. We mine image pairs from the web via (1) grouping images from the same web page and cleaning, (2) annotating metadata for each image with LMMs, and (3) scoring and filtering unqualified image pairs. Eventually, we generate open-ended instructions using LLMs for the remaining image pairs.
@@ 2 Related Work
Composed Image Retrieval. Composed image retrieval (CIR; Vo et al. (2019)) shares the same task form with us. However, all existing benchmarks (Liu et al., 2021; Baldrati et al., 2023; Wu et al., 2021) collect visually similar images first and then write instructions for image pairs. This limits the richness of image relations on these benchmarks and the models developed upon/for them. Recent works on zero-shot CIR either design light-weight modality transformation or adjust training and model to use existing image-caption (Saito et al., 2023; Baldrati et al., 2023; Gu et al., 2024) data. CIReVL (Karthik et al., 2024) uses LLM and LMM on-the-fly for CIR, limiting its efficiency. Please refer to Appendix B for more details of these methods. In terms of constructing training data, CompoDiff (Gu et al., 2023) synthesizes 18M triplets with LLMs and image generative models, following the same pipeline with Brooks et al. (2023). The key difference between their data and ours lies in the image quality and the image relations. As shown in Figure 2, our data comes from natural image pairs found on the same web pages. Thus, our data covers open-ended image relations over a wide distribution, including both visual and beyond-visual ones.
Retrieval with Instruction. Instruction tuning (Ouyang et al., 2022) enables models with strong cross-domain and zero-shot generalization capabilities in retrieving both textual (Su et al., 2023; Asai et al., 2023) and multimodal content (Wei et al., 2023). However, prior efforts focus on unifying different retrieval tasks with manually-written instructions as task prefixes of actual queries, on a hundred-scale basis. In contrast, our approach utilizes million-scale instructions that naturally express user’s search intents.
@@ 3.1 Data Construction for Self-Supervised Training
(2) Metadata Expansion. To provide detailed textual information of images for later LLMs with massive metadata expansion, we annotate images with Alt-texts [^3] https://en.wikipedia.org/wiki/Alt_attribute, image content annotation (ICA) labels [^4] https://cloud.google.com/vision/docs/labels, and captions. We discard images if their Alt-texts are unqualified. For ICA labels, we annotate entities for each image such as general objects and activities. For image captions, we adopt a SOTA LMM — PaLI (Chen et al., 2023a) to generate captions. Each type of metadata provides textual information about images from different perspectives. See Appendix A for more details.
@@ 3.2 MagicLens Model
Model Design. As shown in Figure 4, we adopt a simple dual-encoder architecture with shared parameters and initialize the backbone vision and language encoders with CoCa (Yu et al., 2022) or CLIP (Radford et al., 2021). To enable deep modality integration, we introduce multiple self-attention layers and design a single multi-head attention pooler, compressing the multimodal inputs into a single embedding $\boldsymbol{r}$ for later matching. Additionally, since the retrieval target comprises only an image without accompanying text, we employ an empty text placeholder “” to transform the target into a multimodal input. We denote $\boldsymbol{r_{q}}$ as the embedding for the multimodal query (imageq, text) and $\boldsymbol{r_{t}}$ as the embedding for the target (imaget, “”). Considering the efficiency, we propose MagicLens-B and MagicLens-L, initialized with the base and large checkpoints, respectively.
@@
Model Training. We use a simple contrastive loss to train MagicLens, our model is updated by contrasting the paired query-target against other targets in one training batch. In particular, as the query image itself (imageq) can be a challenging hard negative for the multimodal query (imageq, text), we combine the query image itself and an empty text to encode (imageq, “”) to get $\boldsymbol{r_{t}^{{}^{\prime}}}$ as an additional query negative example. To scale up the number of negative examples, for each query image, we use all query negatives and other target negatives in the same batch. Formally, for the $i$-th training example, the loss function $\mathcal{L}_{i}$ is defined as,
@@ 4.1 Experiment Setup
Composed Image Retrieval. We consider one domain-specific and two open-domain benchmarks to evaluate model’s domain adaptability and its capability on real-world natural images, respectively. FIQ (Wu et al., 2021) is a fashion-domain benchmark with three disjoint retrieval sub-tasks: dress, shirt, and toptee. Following previous work (Saito et al., 2023; Baldrati et al., 2023; Gu et al., 2024), we evaluate on its validation set and report recall averaged over sub-tasks. CIRR (Liu et al., 2021) is the first dataset constructed on natural images (Suhr et al., 2019) with nine pre-defined relations between the query and the target images. It also designs a subset retrieval setting where models retrieve a target image from a dedicated small subset for each query. However, in addition to the limited size of the retrieval pool, it also suffers from false negative issues, as pointed out by Baldrati et al. (2023). We utilize recall (R and Rs) to evaluate standard retrieval and subset retrieval. In contrast, to better align with the real-world large-scale retrieval, CIRCO annotates multiple ground truths for each query and has over 120K natural images (Lin et al., 2014) as the index set. Therefore, we regard CIRCO as our main benchmark. As each query has multiple targets, we adopt mean Average Precision (mAP) as the evaluation metric.
@@
Conditional Image Similarity. GeneCIS (Vaze et al., 2023) is a keyword-conditioned image similarity measurement benchmark. It has four sub-tasks about changing or focusing the attribute or object in the given image. For each query image and keyword, models need to find the most similar images to the query image, conditioned on the given keyword, from a dedicated small subset with 13.8 images on average. For example, in the change-object sub-task with keyword “car” and an image, models need to find another image but with additional cars.
@@ 4.4 Text-to-Image Retrieval
Table 3 shows the comparison between the original encoders and the ones updated after MagicLens training. For text $\rightarrow$ image task, we can observe consistent and non-trivial improvements across all metrics on both datasets. For image $\rightarrow$ text task, we observe marginal drops. These observations show that our training recipe can enhance the backbone encoders for text-to-image retrieval. We can draw the same conclusion with CLIP, which is detailed in Table 17. The improvements might stem from the fact that our multimodality-to-image training task necessitates deep understanding of text instruction, thus improving backbone language encoders. These text-to-image results show that MagicLens can successfully handle three kinds of image retrieval task forms with all strong results, considering the performances on other tasks aforementioned.
@@ 5.1 Data Analysis
Table 5: Performance comparison of CoCa-based MagicLens-B trained with 1M IP2P data against 1M our data. We report averaged R@10 & R@50 on FIQ and averaged R@1 & Rs@1 CIRR for fair comparisons with CompoDiff (Gu et al., 2023).
@@
Data Scaling. To investigate the effect of our data scale on models, we train CoCa-based MagicLens-B on randomly sampled sets of 0.2M, 0.5M, 1M, 5M, 10M, 25M, and the entire 36.7M triplets. Results on five benchmarks and their averaged performance are illustrated in Figure 6. As the data size increases, MagicLens shows enhanced average performance, especially before the 10M mark. This indicates the effectiveness of scaling data.
@@ 5.2 Model Analysis
Model Size vs. Performance. Previous SOTA methods (Gu et al., 2024, 2023) consider using both larger vision and language encoders (Cherti et al., 2023) or using LMMs and LLMs (Karthik et al., 2024) on-the-fly for performance benefits. However, we argue that the model sizes and the correlated efficiencies should also be taken into consideration for real-world deployments. In Figure 7, we visualize the relationship between model size and performance of various models on GeneCIS, CIRCO, and DTIN benchmarks. The results on GeneCIS and CIRCO are from Gu et al. (2024, 2023) and using CLIP-Large, OpenCLIP-Huge, and OpenCLIP-Giant backbones (Radford et al., 2021; Cherti et al., 2023). Results of CIReVL (Karthik et al., 2024) on DTIN and GeneCIS are not fully reported by the authors. We omit the size of ChatGPT used and only count parameters of CIReVL’s other model components (e.g., BLIP2-FLANT5-XXL + OpenCLIP-Giant).
Despite the 50$\times$ smaller size of CoCa-based MagicLens-B (267M) compared to other baselines (e.g., CIReVL with 14.6B), it achieves better performance on these benchmarks, with a significant advantage on the DTIN. This observation demonstrates the high parameter efficiency introduced by the parameter-sharing design in our model and the strong advantage of our data in enabling strong yet small models. Detailed results are in Appendix Table 1415, and 16.
@@
Ablation on Contrastive Loss. Compared to standard contrastive loss, we introduce query images as hard negative examples during training. To investigate the impact of this design, we train CoCa-based MagicLens-B without these hard negatives and report the results in Table 7. As we can see, without query negatives, the performance of MagicLens drops across all benchmarks, significantly on CIRR, CIRCO, and DTIN benchmarks. Also, we find in many cases, this model prefers to rank the query image itself higher than other images during retrieval, regardless of the given instructions. This indicates that differentiating closely similar images is crucial in improving the model’s instruction understanding capabilities. Importantly, although using query negatives seems to limit MagicLens’ ability to find the identical image, the first example in Figure 1 shows MagicLens can generalize to this instruction unseen during training and successfully retrieve the identical image.

### v2

@@ Abstract
Image retrieval, i.e., finding desired images given a reference image, inherently encompasses rich, multi-faceted search intents that are difficult to capture solely using image-based measures. Recent works leverage text instructions to allow users to more freely express their search intents. However, they primarily focus on image pairs that are visually similar and/or can be characterized by a small set of pre-defined relations. The core thesis of this paper is that text instructions can enable retrieving images with richer relations beyond visual similarity. To show this, we introduce MagicLens, a series of self-supervised image retrieval models that support open-ended instructions. MagicLens is built on a key novel insight: image pairs that naturally occur on the same web pages contain a wide range of implicit relations (e.g., inside view of), and we can bring those implicit relations explicit by synthesizing instructions via foundation models. Trained on 36.7M (query image, instruction, target image) triplets with rich semantic relations mined from the web, MagicLens achieves results comparable with or better than prior best on eight benchmarks of various image retrieval tasks, while maintaining high parameter efficiency with a significantly smaller model size. Additional human analyses on a 1.4M-image unseen corpus further demonstrate the diversity of search intents supported by MagicLens. Code and models are publicly available at the Project Website.
@@ 1 Introduction
Figure 1: Top-1 retrieved images using MagicLens and the prior state-of-the-art (SOTA) method (Gu et al., 2024) from a retrieval pool with 1.4M images. The prior SOTA method, while accepting text instructions, primarily retrieves images based on visual similarity to the query image, ignoring the nuances of the text instructions. In contrast, MagicLens excels at retrieving both visually similar images and those that align with the deeper meaning and context of the text instructions — even when the images do not resemble the query. For example, if given a query image of the Burj Al Arab and the instruction “Find other attractions in this country”, it can successfully locate images of the Palm Islands in Dubai.
@@
In this paper, we present MagicLens, a series of self-supervised image retrieval models trained on a wide range of (query image, instruction, target image) triplets that reflect naturally occurring semantic relations, mined from web pages and curated with state-of-the-art (SOTA) foundation models. Specifically, we extract image pairs that naturally occur on the same web page to form positive pairs that carry abundant but natural semantic relations. We then apply both large multimodal models (LMMs; Chen et al. (2023b, a)) and large language models (LLMs; Anil et al. (2023)) to refine the description of such open-ended semantic relation, into open-ended instruction. Figure 2 shows an overview of the data construction pipeline. For example, a camera review website [^1] https://amateurphotographer.com/review/nikon-z5-review/ presenting the image of a Nikon Camera and the image of a Nikon Charger would offer an interesting and non-trivial relation “charger of a product”, which would then be curated by the LMM$+$LLM pipeline, and produce a final instruction find a charger for it. This process produces open-ended instructions that depict diverse semantic relations beyond mere visual similarity, resulting in a large-scale training dataset with 36.7M high-quality triplets over a wide distribution.
With the constructed dataset, we train dual-encoder models called MagicLens, which retrieve images given a query consisting of an image with an instruction. Our models achieve results comparable with or better than prior SOTA methods on eight benchmarks, including various multimodality-to-image and image-to-image retrieval tasks. In addition, MagicLens can retain or even significantly improve the text-to-image retrieval performance of the underlying single-modality encoders. With a 50 times smaller model size than prior SOTA methods, MagicLens outperforms them on multiple benchmarks: CIRCO (Baldrati et al., 2023), Domain Transfer ImageNet (Saito et al., 2023), and GeneCIS (Vaze et al., 2023). To further examine our models’ capabilities in a more realistic scenario, we construct the largest retrieval pool to date with 1.4 million unseen images and perform retrieval given human-written search queries with diverse instructions. The human evaluation finds that MagicLens can successfully satisfy complex and beyond visual search intents, whereas the prior SOTA fails to do so.
Figure 3: Data construction pipeline. We mine image pairs from the web via (1) grouping images from the same web page and cleaning them, (2) annotating metadata for each image with LMMs, and (3) scoring and filtering out unqualified image pairs. Eventually, we generate open-ended instructions using LLMs for the remaining image pairs.
@@ 2 Related Work
Composed Image Retrieval. Composed image retrieval (CIR; Vo et al. (2019)) shares the same task form with us. However, all existing benchmarks (Liu et al., 2021; Baldrati et al., 2023; Wu et al., 2021) collect visually similar images first and then write instructions for image pairs. This limits the richness of image relations on these benchmarks and the models developed upon/for them. Recent works on zero-shot CIR (Saito et al., 2023; Baldrati et al., 2023; Gu et al., 2024) either design light-weight modality transformation or adjust training and model to use existing image-caption data. CIReVL (Karthik et al., 2024) uses LLM and LMM on-the-fly for CIR, limiting its efficiency. Please refer to Appendix B for more details of these methods. In terms of constructing training data, CompoDiff (Gu et al., 2023) synthesizes 18M triplets with LLMs and image generative models, following the same pipeline with Brooks et al. (2023). The key difference between their data and ours lies in the image quality and the image relations. As shown in Figure 2, our data comes from natural image pairs found on the same web pages. Thus, our data covers open-ended image relations over a wide distribution, including both visual and beyond-visual ones.
Retrieval with Instruction. Instruction tuning (Ouyang et al., 2022; Lou et al., 2024) enables models with strong cross-domain and zero-shot generalization capabilities in retrieving both textual (Su et al., 2023; Asai et al., 2023) and multimodal content (Wei et al., 2023). However, prior efforts focus on unifying different retrieval tasks with manually-written instructions as task prefixes of actual queries, on a hundred-scale basis. In contrast, our approach utilizes million-scale instructions that naturally express user’s search intents.
@@ 3.1 Data Construction for Self-Supervised Training
(2) Metadata Expansion. To provide detailed textual information of images for later LLMs with massive metadata expansion, we annotate images with Alt-texts [^3] https://en.wikipedia.org/wiki/Alt_attribute, image content annotation (ICA) labels [^4] https://cloud.google.com/vision/docs/labels, and captions. We discard images if their Alt-texts are unqualified. For ICA labels, we annotate entities for each image, such as general objects and activities. For image captions, we adopt a SOTA LMM — PaLI (Chen et al., 2023a) to generate captions. Each type of metadata provides textual information about images from different perspectives. See Appendix A for more details.
@@ 3.2 MagicLens Model
Model Design. As shown in Figure 4, we adopt a simple dual-encoder architecture with shared parameters and initialize the backbone vision and language encoders with CoCa (Yu et al., 2022) or CLIP (Radford et al., 2021). To enable deep modality integration, we introduce multiple self-attention layers and design a single multi-head attention pooler, compressing the multimodal inputs into a single embedding $\boldsymbol{r}$ for later matching. Additionally, since the retrieval target comprises only an image without accompanying text, we employ an empty text string “” to transform the target into a multimodal input. We denote $\boldsymbol{r_{q}}$ as the embedding for the multimodal query (imageq, text) and $\boldsymbol{r_{t}}$ as the embedding for the target (imaget, “”). Considering the efficiency, we propose MagicLens-B and MagicLens-L, initialized with the base and large checkpoints, respectively.
@@
Model Training. We use a simple contrastive loss to train MagicLens. Our model is updated by contrasting the paired query-target against other targets in one training batch. In particular, as the query image itself (imageq) can be a challenging hard negative for the multimodal query (imageq, text), we combine the query image itself and an empty text to encode (imageq, “”) to get $\boldsymbol{r_{t}^{{}^{\prime}}}$ as an additional query negative example. To scale up the number of negative examples, for each query image, we use all query negatives and other target negatives in the same batch. Formally, for the $i$-th training example, the loss function $\mathcal{L}_{i}$ is defined as,
@@ 4.1 Experiment Setup
Composed Image Retrieval. We consider one domain-specific and two open-domain benchmarks to evaluate the model’s domain adaptability and its capability on real-world natural images, respectively. FIQ (Wu et al., 2021) is a fashion-domain benchmark with three disjoint retrieval sub-tasks: dress, shirt, and toptee. Following previous work (Saito et al., 2023; Baldrati et al., 2023; Gu et al., 2024), we evaluate on its validation set and report recall averaged over sub-tasks. CIRR (Liu et al., 2021) is the first dataset constructed on natural images (Suhr et al., 2019) with nine pre-defined relations between the query and the target images. It also includes a subset retrieval setting where models retrieve a target image from a dedicated small subset for each query. However, in addition to the limited size of the retrieval pool, it also suffers from false negative issues, as pointed out by Baldrati et al. (2023). We utilize recall (R and Rs) to evaluate standard retrieval and subset retrieval. In contrast, to better align with the real-world large-scale retrieval, CIRCO annotates multiple ground truths for each query and has over 120K natural images (Lin et al., 2014) as the index set. Therefore, we regard CIRCO as our main benchmark. As each query has multiple targets, we adopt mean Average Precision (mAP) as the evaluation metric.
@@
Conditional Image Similarity. GeneCIS (Vaze et al., 2023) is a keyword-conditioned image similarity measurement benchmark. It has four sub-tasks about changing or focusing the attribute or object in the given image. For each query image and keyword, models need to find the most similar images to the query image, conditioned on the given keyword, from a dedicated small subset with 13.8 images on average. For example, in the change-object sub-task with keyword “car” and an image, models need to find another image depicting a similar scene but with additional cars.
@@ 4.4 Text-to-Image Retrieval
Table 3 shows the comparison between the original encoders and the ones updated after MagicLens training. For text $\rightarrow$ image task, we can observe consistent and non-trivial improvements across all metrics on both datasets. For image $\rightarrow$ text task, we observe marginal drops. These observations show that our training recipe can enhance the backbone encoders for text-to-image retrieval. We can draw the same conclusion with CLIP, which is detailed in Table 17. The improvements might stem from the fact that our multimodality-to-image training task necessitates deep understanding of text instruction, thus improving backbone language encoders. These text-to-image results, along with the results on image-to-image and multimodality-to-image tasks, show that MagicLens can well handle various forms of image retrieval tasks, all with strong performances.
@@ 5.1 Data Analysis
Table 5: Performance comparison of CoCa-based MagicLens-B trained with 1M IP2P data against 1M our data. We report averaged R@10 & R@50 on FIQ and averaged R@1 & Rs@1 CIRR for comparisons with CompoDiff (Gu et al., 2023).
@@
Data Scaling. To investigate the effect of our data scale on models, we train CoCa-based MagicLens-B on randomly sampled sets of 0.2M, 0.5M, 1M, 5M, 10M, 25M, and the entire 36.7M triplets. Results on five benchmarks and their average performance are illustrated in Figure 6. As the data size increases, MagicLens shows enhanced average performance, especially before the 10M mark. This indicates the effectiveness of scaling data.
@@ 5.2 Model Analysis
Model Size vs. Performance. Previous SOTA methods (Gu et al., 2024, 2023) consider using both larger vision and language encoders (Cherti et al., 2023) or using LMMs and LLMs on-the-fly (Karthik et al., 2024) for performance benefits. However, we argue that the model sizes and the correlated efficiencies should also be taken into consideration for real-world deployments. In Figure 7, we visualize the relationship between model size and performance of various models on GeneCIS, CIRCO, and DTIN benchmarks. The results on GeneCIS and CIRCO are from Gu et al. (2024, 2023) and using CLIP-Large, OpenCLIP-Huge, and OpenCLIP-Giant backbones (Radford et al., 2021; Cherti et al., 2023). Results of CIReVL (Karthik et al., 2024) on DTIN are not fully reported by the authors. We omit the size of ChatGPT used and only count parameters of CIReVL’s other model components (e.g., BLIP2-FLANT5-XXL + OpenCLIP-Giant).
Despite the 50$\times$ smaller size of CoCa-based MagicLens-B (267M) compared to other baselines (e.g., CIReVL with 14.6B), it achieves better performance on these benchmarks, with a significant advantage on the DTIN. This observation demonstrates the high parameter efficiency introduced by the parameter-sharing design in our model and the strong advantage of our data in enabling strong yet small models. Detailed results are in Table 1415, and 16 in Appendix.
@@
Ablation on Contrastive Loss. Compared to standard contrastive loss, we introduce query images as hard negative examples during training. To investigate the impact of this design, we train CoCa-based MagicLens-B without these hard negatives and report the results in Table 7. As we can see, without query negatives, the performance of MagicLens drops across all benchmarks, significantly on the CIRR, CIRCO, and DTIN benchmarks. Also, we find in many cases, this model prefers to rank the query image itself higher than other images during retrieval, regardless of the given instructions. This indicates that differentiating closely similar images is crucial in improving the model’s instruction understanding capabilities. Importantly, although using query negatives seems to limit MagicLens’ ability to find the identical image, the first example in Figure 1 shows MagicLens can generalize to this instruction unseen during training and successfully retrieve the identical image.